OpenStudy (anonymous):

What is the reason for the fourth statement in this proof? definition of collinear points segment addition Subtraction Property of Equality Substitution Property of Equality

OpenStudy (neonumbrella5115):

@qwertyboy56 What proof?

OpenStudy (anonymous):

my bad fergot the picture @neonumbrella5115

OpenStudy (anonymous):

Given: AD = CF BC = DE (Click here to see the image.) Prove: AB = EF Statement Reason 1. AD = CF given 2. AD = AB + BC + CD CF = CD + DE + EF segment addition 3. AB + BC + CD = CD + DE + EF Transitive Property of Equality 4. AB + BC = DE + EF 5. BC = DE given 6. AB = EF Subtraction Property of Equality

OpenStudy (neonumbrella5115):

@qwertyboy56 No problem. So, you need a reason for why AB + BC + CD = CD + DE + EF is equal to AB + BC = DE + EF?

OpenStudy (anonymous):

whats # 4 reason

OpenStudy (anonymous):

4 is the only one missing a reason

OpenStudy (neonumbrella5115):

@qwertyboy56 Yes. I was just making sure. So, what happened to CD in the steps? Where did it go? It cancelled out with the other CD on the other side of the equation. So, what do you do to the CD to make it cancel with the other Cd?

OpenStudy (anonymous):

got it wrong

OpenStudy (anonymous):

the answer is Subtraction Property of Equality